# Enumerative Geometry of Quantum Periods

### v1

@@ Abstract.
We interpret the $q$-refined theta function $\vartheta_{1}$ of a log Calabi-Yau surface $(\mathbb{P},E)$ as a natural $q$-refinement of the open mirror map, defined by quantum periods of mirror curves for outer Aganagic-Vafa branes on the local Calabi-Yau $K_{\mathbb{P}}$. The series coefficients are all-genus logarithmic two-point invariants, directly extending the relation found in [GRZ]. Yet we find an explicit discrepancy at higher genus in the relation to open Gromov-Witten invariants of the Aganagic-Vafa brane. Using a degeneration argument, we express the difference in terms of relative invariants of an elliptic curve. With $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ the toric blow up of a point, we use the Topological Vertex [AKMV] to show a correspondence between open invariants of $K_{\mathbb{P}}$ and closed invariants of $K_{\widehat{\mathbb{P}}}$ generalizing a variant of [CCLT, LLW] to arbitrary genus and winding. We also equate winding-1, open-BPS invariants with closed Gopakumar-Vafa invariants.
@@ 1. Introduction
In [GRZ], the authors found new enumerative interpretations involving periods of the mirror curve of a toric Calabi-Yau threefold, $Y$111 Other enumerative interpretations had previously been found [CLL, CLT, CCLT, CLLT, LLW]. See in particular Theorem 1.1 and Example 4.6 of [LLW] and Example 6.5 (3) of [CCLT]. The conjecture in [GRZ] was partly proved in [You]. The Picard-Fuchs equations determine the periods, and in particular the open mirror map, and its coefficients have two equal interpretations: (1) two-point relative invariants used to define the first theta function, or proper Landau-Ginzburg potential; (2) open invariants of an Aganagic-Vafa brane on $Y.$ In this paper, we extend (1) to higher genus by relating the $q$-refined quantum periods of the quantum mirror curve. We find a discrepancy in (2) at higher genus, and give an explicit expression for the correction by analyzing the degeneration formula. Here, by a $q$-refinement we mean the following.
@@
which $q$-refines $-\frac{1}{3}F(-z)=-2z-15z^{2}-\cdots.$ The non-refined period $-\frac{1}{3}F(-z)$ leads to open GW predictions for the AV brane in framing zero, which for genus zero and winding one in low degrees gives $2,5,32,\cdots$. These same numbers were also found in [GRZ] to be two-pointed logarithmic invariants in $\mathbb{P}^{2}$ in genus zero. The same procedure with the quantum period, gives a $q$-refinement of these numbers — specifically
@@
— see Example 2.27. However, the $q$-dependence of the OGW invariants for the AV brane is computed via the topological vertex [AKMV], Equation (9.10) to be
@@
The aim of this paper is first to prove that the quantum periods computed via the mirror symmetry procedure determine the all-genus logarithmic two-point invariants, then to explain the discrepancy from the OGW invariants.
1.1. Main statements
Here we describe our key findings. The combinatorial Laurent series methods used in §4 work in any dimension, but for the treatment of higher genus invariants with $\lambda$-insertions, we have to restrict to surfaces.
Let $\mathbb{P}$ be a smooth projective Fano surface and let $E$ be a smooth anticanonical divisor, so $(\mathbb{P},E)$ is a log Calabi-Yau pair. Section 1.2 contains our full notation conventions and Section 2 relevant definitions.
1.1.1. Quantum periods
@@
where the wavefunction $\Psi(X)$ is a solution to the difference equation $p\cdot\Psi=0$ defined by the quantum mirror curve, see §2.5.
@@
To compute $\underline{a}(z,q)$ from [ACDKV], one can write $Y(X):=\Psi(qX)/\Psi(X)$ as an ansatz $Y(X)=\sum_{k\geq 0}Y_{k}(X)s^{k}$, with $s=q^{\frac{1}{2}}z$ and then solve order-by-order in $s$, see §2.5. Theorem 2.23 states that one can work with the alternative definition $a(z,q)$, which can be computed combinatorially.
@@
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the Chern classes of the Hodge bundle.
@@
1.1.2. Natural $q$-refinements of mirror maps
@@
$M(Q,q)=1+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}\frac{1}{\beta\cdot{E}-1}R_{g,(1,\beta\cdot{E}-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$
@@
Theorem 4.2 and Corollary 4.3 extend the results of [GRZ] to the $q$-refined setting. We prove Theorem 4.2 using combinatorial power series results that are proved in §A. These include a Lagrange inversion formula for formal Laurent series and an exponentiation formula for classical periods of these. The latter follows from a relation for partial ordinary Bell polynomials proved in [BGW]. Then Corollary 4.3 follows from the tropical correspondence results in [GRZ]. The combinatorial arguments of §4 work in any dimension. For the q-refined tropical correspondence of §3 we need surfaces.
1.1.3. Open-log correspondence and correction terms
In Section 5, we prove an all genera open-log correspondence between two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and certain open invariants of $K_{\mathbb{P}}.$ Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up at a point with exceptional curve $C$, and let $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ be the genus-$g$ Gopakumar-Vafa invariant of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-C$ defined by multiple cover formulas [GV1] [GV2]. We first show the following theorem relating $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.
@@ 1.1.3. Open-log correspondence and correction terms
We prove Theorem 5.2 by using higher genus analogues of arguments in Theorem 1.1, [GRZ]. The discrepancy term $\Delta$ comes from application of the $g\geq 0$ log-local principle of [BFGW] and a blow up formula for log invariants in Corollary 6.6, [GRZ], and is expressed by the stationary Gromov-Witten theory of $E$ and, after fixing a genus-$g$, the two-pointed log invariants $R_{h,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ for $h<g.$
1.1.4. Winding-w open/closed correspondence
In Section 5, we additionally assume that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\to\mathbb{P}$ is a toric blow up, so that $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. Let $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the genus-$g$, winding-$w$, open-BPS invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ in framing-0, curve class $\beta+w\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$, and in winding profile $\vec{k}$ defined by multiple cover formulas for open Gromov-Witten invariants [MV] (see Section 5 for more details). In Theorem 5.3, we prove an equality between winding-$w$ open partition functions of $K_{\mathbb{P}}$ and a certain closed partition function of $K_{\widehat{\mathbb{P}}}$. Specializing Theorem 5.3 to winding-1, open invariants gives the following correspondence of enumerative invariants.
Theorem (Corollary 5.4).
Suppose $\mathbb{P}$ is a toric Fano surface, $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point with exceptional curve $C$, and let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. Then we have,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$
In genus-0, Corollary 5.4 is equivalent to Theorem 1.1, [LLW]. In Section 5.6, we provide a computational verification for Corollary 5.4 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$ in low degrees and all genera.
By substituting Corollary 5.4 into Theorem 5.2, we show an all genera open-log correspondence in Theorem 5.8 relating two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ for smooth log Calabi-Yau pairs. In genus-00, Theorem 5.8 reduces to the open-log correspondence established in [GRZ]. We give an explicit example of Theorem 5.8 in Example 5.10.
@@ 1.2. Notation
The mirror curve is a moduli space for A-branes on the local Calabi-Yau threefold $K_{\mathbb{P}}$, see §2.2. We denote a local coordinate on the mirror curve by $x$. Mirror dual to $x$ is a symplectic area modulus of B-branes, which we denote by $U$.
@@
We write $R_{g,(p,q)}(\mathbb{P}(\log E),\beta)$ for the logarithmic Gromov-Witten invariant counting curves on $\mathbb{P}$ of genus-$g$ and class $\beta$ intersecting the divisor $E$ in two points, a prescribed point with multiplicity $p$ and a non-prescribed point with multiplicity $q$. Let $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$ be a disc class with boundary on an outer Aganagic-Vafa brane $L$. We write $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ for the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0 and winding profile $\vec{k}$, and $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be its corresponding open-BPS invariant. We write $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be the genus-$g$ local Gromov Witten invariant of $\widehat{\mathbb{P}}$ in curve class $\pi^{*}\beta-wC$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be its corresponding Gopakumar-Vafa invariant.
@@ 1.3. Acknowledgments
We would like to thank Kwokwai Chan, Melissa Liu, Daniel Krefl, Ezra Getzler and Jörg Teschner for helpful exchanges. T.G. would like to thank the MFO Oberwolfach and the organizers of the workshop Real and Logarithmic Enumerative Geometry for the opportunity to present the results of this project. E.Z. has been supported by NSF grant DMS-2104087. H.R. is supported by the NFR Fripro grant Shape2030.
@@ 2.1. Fano surfaces
Let $\mathbb{P}$ be a smooth surface with ample anticanonical bundle, i.e., one of the $8$ del Pezzo surfaces given by $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or the blow up of $\mathbb{P}^{2}$ in $0\leq k\leq 6$ points. Let $\mathbb{P}_{\Sigma^{\prime}}$ be a Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. If $\mathbb{P}$ is toric, we can simply take $\mathbb{P}_{\Sigma^{\prime}}=\mathbb{P}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the primitive integral ray generators of $\Sigma^{\prime}$. The ray generators are exactly the vertices of $\Delta^{\star}$. Any mutation of $\Delta^{\star}$ gives another Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration [Ilt]. Let $\Sigma$ be the subdivision of $\Sigma^{\prime}$ whose ray generators are all integral points on the boundary of $\Delta^{\star}$, not just the vertices. Then $\mathbb{P}_{\Sigma}$ is a smooth semi-Fano toric surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. Again, if $\mathbb{P}$ is toric we can take $\mathbb{P}_{\Sigma}=\mathbb{P}$. There are exactly 16 possibilities for $\mathbb{P}_{\Sigma^{\prime}}$, hence for $\mathbb{P}_{\Sigma}$, corresponding to the 16 reflexive polygons. See §3.2 and [CPS] for more details and [Gra1], Figure 1.3, for a complete list of cases.

### v2

@@ Abstract.
We interpret the $q$-refined theta function $\vartheta_{1}$ of a log Calabi-Yau surface $(\mathbb{P},E)$ as a natural $q$-refinement of the open mirror map, defined by quantum periods of mirror curves for outer Aganagic-Vafa branes on the local Calabi-Yau $K_{\mathbb{P}}$. The series coefficients are all-genus logarithmic two-point invariants, directly extending the genus-zero relation found in [GRZ]. We show that the winding-$1$ open-closed correspondence of Chan and Lau-Leung-Wu [Cha, LLW] generalizes to higher genus in the toric situation by using the Topological Vertex. We conjecture a more general, higher-winding open-closed relation involving LMOV invariants. Yet we find an explicit discrepancy with the log invariants of $(\mathbb{P},E)$, expressible in terms of stationary Gromov-Witten invariants of an elliptic curve.
@@ 1. Introduction
In [GRZ], the first three authors found new enumerative interpretations involving periods of the mirror curve of a toric Calabi-Yau threefold, $Y$111 Other enumerative interpretations had previously been found [CLL, CLT, CCLT, CLLT, LLW]. See in particular Theorem 1.1 and Example 4.6 of [LLW] and Example 6.5 (3) of [CCLT]. The conjecture in [GRZ] was partly proved in [You]. The Picard-Fuchs equations determine the periods, and in particular the open mirror map, and its coefficients have two equal interpretations: (1) two-point relative invariants used to define the first theta function, or proper Landau-Ginzburg potential; (2) open invariants of an Aganagic-Vafa brane on $Y.$ In this paper, we extend (1) to higher genus by relating the $q$-refined quantum periods of the quantum mirror curve. We find a discrepancy in (2) at higher genus, and give an explicit expression for the correction by analyzing the degeneration formula. Here, by a $q$-refinement we mean the following.
@@
which $q$-refines $-\frac{1}{3}F(-z)=-2z-15z^{2}-\cdots.$ The non-refined period $-\frac{1}{3}F(-z)$ leads to open Gromov-Witten predictions for the Aganagic-Vafa (AV) brane in framing zero, which for genus zero and winding one in low degrees gives $2,5,32,\cdots$. These same numbers were also found in [GRZ] to be two-pointed logarithmic invariants in $\mathbb{P}^{2}$ in genus zero. The same procedure with the quantum period, gives a $q$-refinement of these numbers — specifically
@@
— see Example 2.28. However, the $q$-dependence of the open Gromov-Witten invariants for the AV brane is computed via the topological vertex [AKMV], Equation (9.10) to be
@@
The aim of this paper is first to prove that the quantum periods computed via the mirror symmetry procedure determine the all-genus logarithmic two-point invariants, then to explain the discrepancy from the open Gromov-Witten invariants.
We now describe our results. Let $\mathbb{P}$ be a smooth projective Fano surface [^2] The combinatorial Laurent series methods used in Section 4 work in any dimension, but for the treatment of higher genus invariants with $\lambda$-insertions, we have to restrict to surfaces. and let $E$ be a smooth anticanonical divisor, so $(\mathbb{P},E)$ is a log Calabi-Yau pair. Section 1.4 contains our full notation conventions and Section 2 relevant definitions.
1.1. Quantum periods
@@
where the wavefunction $\Psi(X)$ is a solution to the difference equation $p\cdot\Psi=0$ defined by the quantum mirror curve, see Section 2.5.
@@
To compute $\underline{a}(z,q)$ from [ACDKV], one can write $Y(X):=\Psi(qX)/\Psi(X)$ as an ansatz $Y(X)=\sum_{k\geq 0}Y_{k}(X)s^{k}$, with $s=q^{\frac{1}{2}}z$ and then solve order-by-order in $s$, see Section 2.5. Theorem 2.24 states that one can work with the alternative definition $a(z,q)$, which can be computed combinatorially.
@@
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the top Chern classes of the Hodge bundle.
@@
1.2. Natural $q$-refinements of mirror maps
@@
$M(Q,q)=1+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}\frac{(-1)^{\beta\cdot E}}{\beta\cdot{E}-1}R_{g,(1,\beta\cdot{E}-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$
@@
Theorem 4.2 and Corollary 4.3 extend the results of [GRZ] to the $q$-refined setting. We prove Theorem 4.2 using combinatorial results that are proved in Section B. These include a Lagrange inversion formula for formal Laurent series and an exponentiation formula for classical periods of these. The latter follows from a relation for partial ordinary Bell polynomials proved in [BGW]. Then Corollary 4.3 follows from the tropical correspondence results in [GRZ]. The combinatorial arguments of Section 4 work in any dimension. For the $q$-refined tropical correspondence of Section 3 we need surfaces.
1.3. Open-log correspondence
In Section 5, we prove an all-genus, open-log correspondence between two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and certain open invariants of $K_{\mathbb{P}}.$
1.3.1. Log-local correspondence and correction terms
Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up at a point with exceptional curve $C$, and let $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ be the genus-$g$ Gopakumar-Vafa invariant of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-C$ defined by multiple cover formulas [GV1] [GV2]. In Section 5.2, we first show the following theorem relating $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.
@@ 1.3.1. Log-local correspondence and correction terms
We prove Theorem 5.2 by using higher genus analogues of arguments in [GRZ], Theorem 1.1. The discrepancy term $\Delta$ comes from application of the $g\geq 0$ log-local principle of [BFGW] and a blow up formula for log invariants in [GRZ], Corollary 6.6 and is expressed by the stationary Gromov-Witten theory of $E$ and, after fixing a genus-$g$, the two-pointed log invariants $R_{h,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ for $h<g.$
1.3.2. Open-closed correspondence
In Section 5.3, we additionally assume that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\to\mathbb{P}$ is a toric blow up, so that $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. Let $L\subset K_{\mathbb{P}}$ be an outer Aganagic-Vafa brane. Denote $N^{\mathrm{LMOV}}_{g,\nu}(K_{\mathbb{P}}/L,\beta)$ to be the genus-$g$, LMOV invariant of $L$ in representation $\nu$, and $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ be the genus-$g$, winding-$w$, open-BPS invariant of $L$ in winding profile $\vec{k}$, curve class $\beta+w\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$ defined by multiple cover formulas [MV] (see beginning of Section 5 and Section 5.3 for more details of these two invariants). The LMOV invariants $N^{\mathrm{LMOV}}_{g,\nu}$ are related to the open-BPS invariants $n^{open}_{g}$ by a linear transformation (Equation 5.8). In Theorem 5.3, for fixed $w\in\mathbb{Z}_{>0}$, we prove an equality relating LMOV invariants of $L\subset K_{\mathbb{P}}$ in representations of length $w$ to the closed partition function of $K_{\widehat{\mathbb{P}}}$ in curve classes $\pi^{*}\beta-wC$. Specializing Theorem 5.3 to $w=1,2$ gives the following correspondence of enumerative invariants.
Theorem (Corollaries 5.5, 5.7).
Suppose $\mathbb{P}$ is a toric Fano surface, $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point with exceptional curve $C$, and let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. Then for $\beta\in NE(\mathbb{P}),g\geq 0,w=1,2$, we have,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)=(-1)^{g}N^{LMOV}_{g,(w)}(K_{\mathbb{P}}/L,\beta)$
where $(w)$ denotes the representation given by a Young Tableau of a single row of length $w$.
In genus-0, Corollary 5.5 is equivalent to [LLW], Theorem 1.1 after converting from LMOV invariants to open Gromov-Witten invariants using Equation 5.8 and the open multiple cover formula [MV]. In Section 5.5, we provide a computational verification for Corollary 5.5 for $\mathbb{P}=\mathbb{P}^{2}$ in low degrees and all-genus. That the theorem holds for all positive windings $w$ is the content of Conjecture 5.8. We provide partial verification with LMOV and Gopakumar-Vafa invariants listed in Appendix A.
By substituting Corollary 5.5 into Theorem 5.2, we show an all-genus open-log correspondence in Theorem 5.13 relating two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ for smooth log Calabi-Yau pairs. In genus-00, Theorem 5.13 reduces to the open-log correspondence established in [GRZ]. We give an explicit example of Theorem 5.13 in Example 5.14.
@@ 1.4. Notation
The mirror curve is a moduli space for A-branes on the local Calabi-Yau threefold $K_{\mathbb{P}}$, see Section 2.2. We denote a local coordinate on the mirror curve by $x$. Mirror dual to $x$ is a symplectic area modulus of B-branes, which we denote by $U$.
@@
We write $R_{g,(p,q)}(\mathbb{P}(\log E),\beta)$ for the logarithmic Gromov-Witten invariant counting curves on $\mathbb{P}$ of genus-$g$ and class $\beta$ intersecting the divisor $E$ in two points, a prescribed point with multiplicity $p$ and a non-prescribed point with multiplicity $q$. Let $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$ be a disc class with boundary on an outer Aganagic-Vafa brane $L$. We write $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ for the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0 and winding profile $\vec{k}$, and $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be its corresponding open-BPS invariant. We write $N^{LMOV}_{g,\nu}(K_{\mathbb{P}}/L,\beta)$ to be the genus-$g$, LMOV invariant of $L\subset K_{\mathbb{P}}$ in representation $\nu$. We write $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be the genus-$g$ local Gromov Witten invariant of $\widehat{\mathbb{P}}$ in curve class $\pi^{*}\beta-wC$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be its corresponding Gopakumar-Vafa invariant.
@@ 1.5. Acknowledgments
We would like to thank Kwokwai Chan, Babak Haghighat, Melissa Liu, Daniel Krefl, Ezra Getzler and Jörg Teschner for helpful exchanges. E.Z. has been supported by NSF grants DMS-2104087 and DMS-2505652. H.R. is supported by the NFR Fripro grant Shape2030.
@@ 2.1. Fano surfaces
Let $\mathbb{P}$ be a smooth surface with very ample anticanonical bundle, i.e., one of the $8$ del Pezzo surfaces given by $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or the blow up of $\mathbb{P}^{2}$ in $0\leq k\leq 6$ points. Let $\mathbb{P}_{\Sigma^{\prime}}$ be a Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. If $\mathbb{P}$ is toric, we can simply take $\mathbb{P}_{\Sigma^{\prime}}=\mathbb{P}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the primitive integral ray generators of $\Sigma^{\prime}$. The ray generators are exactly the vertices of $\Delta^{\star}$. Any mutation of $\Delta^{\star}$ gives another Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration [Ilt]. Let $\Sigma$ be the subdivision of $\Sigma^{\prime}$ whose ray generators are all integral points on the boundary of $\Delta^{\star}$, not just the vertices. Then $\mathbb{P}_{\Sigma}$ is a smooth semi-Fano toric surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. Again, if $\mathbb{P}$ is toric we can take $\mathbb{P}_{\Sigma}=\mathbb{P}$. There are exactly 16 possibilities for $\mathbb{P}_{\Sigma^{\prime}}$, hence for $\mathbb{P}_{\Sigma}$, corresponding to the 16 reflexive polygons. See Section 3.2 and [CPS] for more details and [Gra1], Figure 1.3, for a complete list of cases.